A small drop of water is suspended motionless in air by a unilform electric field that is directed upward and has a a magnitude of 8560 N/C. The mass of the water drop is 3.64 x 10ko. (a) Is the excess charge on the water drop positive or negative? O positive O negative This answer has not been graded yo (b) How many excess electrons or protons reside on the drop? Additional Materials Lu Section 18
Answer #1

As water drop is suspended motionless in air, electric force on it must be in upward direction.

Force in positive charge is in the direction of electric field and that on negative charge is in direction, opposite to that of electric field.

Electric field and electric force on water drop is upward so excess charge on water drop is positive.

electric force (Eq) = weight of drop (mg)

q = mg/E

Number of excess protons(n)= q/ charge of proton (e)

n = mg / Ee = 3.84×10^-9 × 9.81/8560 × 1.6 ×10^-19

= 2.7 ×10^7